OpenStudy (abbycross167):

Write the point-slope (y - k = m(x - h)) form of the equation from the information given. The line goes through the points (6, 5) and (9, 6) Can someone please help me?

OpenStudy (pawanyadav):

Put above points in eq. y=mx + c you will get m and c

OpenStudy (abbycross167):

where would the points go in the equation? @Pawanyadav

OpenStudy (johnweldon1993):

So we have \(\large y - k = m(x - h)\) We first need to get the slope of the line 'm' in that equation \[\large m = \frac{y_2 - y_1}{x_2 - x_1}\] So what would the slope be?

OpenStudy (abbycross167):

6/9-5/6?

OpenStudy (pawanyadav):

y=mx +c 5=6m+c. first equation Again, 6=9m+c. 2nd equation 2 equation and 2 variable, I think you can solve these equations

OpenStudy (johnweldon1993):

Not quite...just plug it in the formula up there as I have written it \[\large m = \frac{6-5}{9-6} = \frac{1}{3}\] So we know the slope 'm' = 1/3 right?

OpenStudy (abbycross167):

Yes sir

OpenStudy (johnweldon1993):

So now that we have 'm' lets rewrite out point slope equation \[\large y - k = \frac{1}{3}(x - h)\] Now we simply need to pick a point *either of your points, doesn't matter which one* and plug it in In your question here, the points are in the form \(\large (h,k)\) so whatever point you pick, the first number will be your 'h' value and the second will be your 'k' value

OpenStudy (abbycross167):

so it'll look like y-5=1/3(x-6)?

OpenStudy (johnweldon1993):

Exactly!

OpenStudy (abbycross167):

Oh ok!! Awesome thank you soo much!!! You're a great help

OpenStudy (johnweldon1993):

No problem :)